# Geometric Bounds for Persistence

### v1

@@ 6 Final Remarks
While our results providing bounds on lifespans via widths are of theoretical nature, they can be quite useful in applications in the following way. Computing the exact Alexandrov, Urysohn, or Kolmogorov $k$-width for a given set $\mathcal{X}$ can be computationally challenging, but note that by definition, these widths are basically infima over some measurements related to a $k$-dimensional space. While finding this optimal $k$-dimensional space is very hard, any relevant measure taken over a meaningful and well-chosen $k$-dimensional space gives an upper bound for these lifespans. Therefore, while computing the optimal such bound may not be feasible, these results can be effectively used to give some rough upper bounds for the lifespans of relevant topological features.
We would like to add that while we bring ideas from metric geometry to applied algebraic topology, our results can be used in the reverse direction, too. In particular, the lifespans are upper bounded by various width notions in Section 3, which implies that the lifespans gives lower bounds for the width estimation. In other words, for a given metric space $\mathcal{X}$, the maximum lifespan of a class in $\mathrm{Spec}_{k}(\mathcal{V}_{\bullet}(\mathcal{X}))$ gives a lower bound for the relevant $k$-width (e.g., $\operatorname{AW}_{k}(\mathcal{X}),\operatorname{UW}_{k}(\mathcal{X})$), which can be used as an effective computational approach to estimate $k$-widths.
Original PCA vs. Lifespans.
@@
In Section 3.5, we give bounds for Čech lifespans via the variances $\{\nu_{k}\}$ induced by ${}_{\infty}$ (Section 1) As one can easily notice, in order to have such a rigorous bound, we modified the usual PCA definition and considered an $\ell^{\infty}$-variant. However, this makes the ${}_{\infty}$ structure highly sensitive to outliers, in contrast with the original PCA (with $\ell^{2}$-norm). On the other hand, standard PCA is a mainstream, highly effective dimension reduction tool for real-life applications with several computational techniques available. While our results do not say anything about the relation between the original PCA and PH, an experimental result relating the lifespans of bars in the persistence diagram of the Čech filtration $\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbb{R}^{N})$ with variances $\widetilde{\nu}_{k+1}(\mathcal{X})$ of original PCA would be very interesting, for a given finite set of points $\mathcal{X}\subset\mathbb{R}^{N}$. It would particularly interesting and useful for real like applications to to carry out a statistical comparative analysis for random finite subsets $\mathcal{X}$ in $\mathbb{R}^{N}$ (and for $k\geq 1$).
@@
One can notice that when describing the width based arguments in Section 3, we first introduce a $k$-dimensional optimal core $\Gamma_{k}$ for a given set $\mathcal{X}$, then the $k$-width $\mathrm{W}_{k}(\mathcal{X})$ is defined as some kind of “distance” from $\mathcal{X}$ to $k$-core $\Gamma_{k}$. Hence, when $\mathrm{W}_{k}(\mathcal{X})$ is small, in metric geometry, $\mathcal{X}$ is considered as “essentially $k$-dimensional”. There is a similar notion in statistics called Principal Curves and Surfaces [HS89, Del01, OE11]. While principal curves and surfaces are defined as $1$- and $2$-dimensional objects, one can easily generalize the idea to any dimension $k$, e.g. principal $k$-manifolds. In our setting, for a compact subset $\mathcal{X}$ in a Banach space $\mathbf{Y}$, principal curves and surfaces can be considered as $k$-dimensional objects $\Sigma_{k}$ which minimizes $\ell^{2}$-distance from $\mathcal{X}$ to $\Sigma_{k}$ for $k=1,2$ with some normalization condition on $\Sigma_{k}$. In this respect, one can see principal curves and surfaces as a nonlinear generalization of principal component analysis (PCA).
In that respect, for a given set $\mathcal{X}$, our $k$-cores and principal $k$-manifolds can be considered as analogous objects. In Section 3.5, we defined ${}_{\infty}$ as $\ell^{\infty}$-version of the original $\ell^{2}$ PCA . Similarly, our $k$-cores $\Gamma_{k}$ minimize the $\ell^{\infty}$-distance between $\mathcal{X}$ and $\Gamma_{k}$ while principal $k$-manifolds $\Sigma_{k}$ minimizes the $\ell^{2}$-distance between $\mathcal{X}$ and $\Sigma_{k}$. By using this analogy, as principal curves and surfaces are suggested as dimension reduction method, one can consider our $k$-cores represent the essential structure of $\mathcal{X}$ when $\mathrm{W}_{k}(\mathcal{X})$ is small. Furthermore, just like the previous part (PCA vs ${}_{\infty}$), it would be interesting to make a statistical study of the relation between the Čech lifespans and $\ell^{2}$-distance to principal $k$-manifolds.
@@ Appendix A Auxiliary Properties of Widths
Suppose $\operatorname{UW}_{k}(\mathcal{X})<\delta$, and let us show that $\operatorname{AW}_{k}(\mathcal{X}\subset\mathbf{Y})<\delta$. There is a continuous map $f:\mathcal{X}\to\Delta^{k}$ to a $k$-dimensional complex with fibers of diameter $<\delta$. Subdivide $\Delta^{k}$ very finely, so that the preimage of any open star of $\Delta^{k}$ under the map $f$ has diameter $<\delta$ [^12] Recall that the open star $S_{v}$ of a vertex $v\in\Delta^{k}$ is the union of the relative interiors of all simplices of $\Delta^{k}$ that contain $v$. For each vertex $v\in\Delta^{k}$, pick a point $c_{v}$ in the preimage of the open star $S_{v}$ (unless this preimage is empty; in this case, we can safely remove $S_{v}$ from $\Delta^{k}$). The ball $B_{\delta}(c_{v})\subset\mathbf{Y}$ of radius $\delta$ centered at $c_{v}$ covers $f^{-1}(S_{v})$. Consider an auxiliary map $\gamma:\Delta^{k}\to\mathbf{Y}$ defined by sending any vertex $v\in\Delta^{k}$ to $c_{v}\in\mathbf{Y}$, and then extending linearly on $\Delta^{k}$. Compose this map with $f$, and consider $\gamma\circ f:\mathcal{X}\to\mathbf{Y}$. Its image is a simplicial complex of dimension at most $k$, and to complete the proof it suffices to show that every point $x\in\mathcal{X}$ is moved by distance $\|x-\gamma(f(x))\|<\delta$. Let $f(x)$ lie in the relative interior of a simplex of $\Delta^{k}$ with the vertices $v_{0},\ldots,v_{m}$. Then $x\in f^{-1}(S_{v_{i}})\subset B_{\delta}(c_{v_{i}})$, and $\|x-c_{v_{i}}\|<\delta$, where $0\leq i\leq m$. By construction of $\gamma$, the point $\phi(f(x))$ lies in the convex hull of the points $c_{v_{i}}$, $0\leq i\leq m$. Therefore, $\|x-\gamma(f(x))\|$ does not exceed the maximum of $\|x-c_{v_{i}}\|$ over $0\leq i\leq m$, and this maximum is less than $\delta$. ∎
@@
Denote $\mathcal{X}=\bigcap\mathcal{X}_{i}$, $w=\operatorname{TW}_{k}(\mathcal{X}\subset\mathbf{Y})$, and let $f:\mathcal{X}\to\mathbf{Y}$ be a witness map: a continuous map whose image lies in a finite contractible $k$-dimensional simplicial complex $\Delta$ and such that $\|x-f(x)\|<w+\epsilon/3$ for all $x\in\mathcal{X}$ and some arbitrarily chosen $\epsilon>0$. For each point $y\in\Delta$, the fiber $f^{-1}(y)$ lies in the open ball $B_{w+\epsilon/3}(y)$. Since $f$ is continuous on the compact set $\mathcal{X}$, it is uniformly continuous, hence there is a tiny radius $\rho>0$ such that such that the “thickened fiber” $f^{-1}(B_{\rho}(y))$ lies in the open ball $B_{w+2\epsilon/3}(y)$, for each $y\in\Delta$. Since $\Delta$ is a finite complex of dimension $k$, there is a finite closed cover $\bigcup D_{j}=\Delta$ such that its multiplicity is $k+1$, its nerve is homotopy equivalent to $\Delta$, each $D_{j}$ fits in a ball $B_{\rho}(y_{j})$, and $C_{j}:=f^{-1}(D_{j})$ fits in the ball $B_{w+2\epsilon/3}(y_{j})$.
@@
is attained and positive. Take a positive $\delta$ smaller than $\epsilon/3$ and also smaller than each $\delta_{J}$ over all collections $J$ such that $\bigcap\limits_{j\in J}C_{j}=\varnothing$. Consider the open cover $\{U_{j}\}$ of $\mathcal{X}$, where $U_{j}=\mathcal{N}_{\epsilon}(C_{j}\subset\mathbf{Y})$. It has the same nerve as $\{C_{i}\}$, and each $U_{j}$ is contained in $B_{w+\epsilon}(y_{j})$. We have
$\mathcal{X}\subset\bigcup_{j}U_{j}\subset\mathcal{N}_{\epsilon}(\mathcal{X}\subset\mathbf{Y}).$
Now pick a partition of unity $\{\psi_{j}\}$ subordinate to $\{U_{j}\}$. Use it to map $\bigcup_{j}U_{j}$ to its nerve, which has the same dimension and homotopy type as $\Delta$. Namely, a point $x\in\bigcup_{j}U_{j}$ is mapped to $\sum\limits_{j}\psi_{j}(x)y_{j}$. Since $U_{j}\subset B_{w+\epsilon}(y_{j})$, it follows the distance from $x$ to its image is less than $w+\epsilon$.
@@
Here we only work in $\mathbf{Y}=\mathbb{R}^{N}$ with Euclidean metric, and we are interested in 00-übercontractible sets, which we call übercontractible. A rich source of übercontractible sets comes from the cut-locus construction.
@@
The cut-locus $C^{\partial K}$ of the boundary of a convex body $K$ is übercontractible as long as it satisfies the following “tameness” assumption: $C^{\partial K}$ consists of finitely many compact convex sets.
The tameness assumption is introduced here in order to apply a lemma from Morse theory for the distance functional (B.3 below). It is highly likely that both the lemma and theorem can be stated and proven in greater generality, removing the tameness assumption completely. For now, we just remark that tame convex bodies in the sense of B.2 are $C^{2}$-dense among all convex bodies.
The key lemma in the proof is the following application of Morse theory for the distance functional. It was applied in [BHK15, Section 3] to give a sufficient condition for a set to be contractible, but the proof there, in fact, guarantees that the set is übercontractible.
Lemma B.3 (cf. [BHK15, Theorem 3.2]).
Let $\mathcal{T}\subset\mathbf{Y}$ be a union of finitely many compact convex sets. For each $x\notin\mathcal{T}$ consider the largest ball centered at $x$ whose interior does not meet $\mathcal{T}$. Let $B^{\mathcal{T}}(x)$ be the set of the points where this ball touches $\mathcal{T}$. Suppose that for each $x\notin\mathrm{conv}(B^{\mathcal{T}}(x))$. Then $\mathcal{T}$ is übercontractible.
Proof of B.2 for polytopes.
@@
Unfortunately, the case of the general convex body $K$ cannot be proven by approximating $K$ with polytopes $K_{i}$, because in order to guarantee $C^{0}$-convergence $C^{\partial K_{i}}\to C^{\partial K}$, we cannot get away with just polytopes. We sketch an argument that works for fairly general convex bodies, modulo some technicalities.
Sketch of the proof of B.2 for more general convex bodies, including tame ones.
Suppose our convex body $K$ can be approximated by a convex body $K^{\prime}$, satisfying the following three properties (these will be sufficient to prove the übercontractibility of $C^{\partial K}$).
@@
    The boundary of $K^{\prime}$ is $C^{2}$-smooth and strongly convex in the sense that its second fundamental form is positive definite.
@@
    The cut-locus of $\partial K^{\prime}$ consists of finitely many compact convex sets (we assume this to be able to use B.3).
  3. 3.
    The cut-locus of $\partial K^{\prime}$ approximates the cut-locus of $\partial K$ well enough so that the übercontractibility of the former implies the übercontractibility of the latter.
The third property is somewhat vague, so we start by arguing that the tameness assumption suffices. If $K$ is tame, we can inflate it slightly to get a new convex body $K^{\prime}$ that satisfies the first property, while preserving cut-locus $C^{\partial K^{\prime}}=C^{\partial K}=\mathcal{T}$ (which already satisfies the second property by the tameness assumption). This amounts to choosing $r^{\partial K^{\prime}}(x)$ for $x\in\mathcal{T}$ carefully, so that on every face of $\mathcal{T}$, $r^{\partial K^{\prime}}$ is $C^{1}$-close to $r^{\partial K}$, and $r^{\partial K^{\prime}}$ is strongly convex. The details are omitted.
In the rest of the proof we assume that $K$ satisfies the first two properties above, and we will verify the assumption of B.3 for $\mathcal{T}=C^{\partial K}$ to show that it is übercontractible. Pick points $x\in K\setminus\mathcal{T}$, and $y\in B^{\mathcal{T}}(x)$. Our goal is to prove that $\langle y-x,n\rangle>0$, where $n$ is the inner normal to $\partial K$ at $b^{\partial K}(x)$. This will imply that $x\notin\mathrm{conv}(B^{\mathcal{T}}(x))\subset\{z\in Y:\langle z-x,n\rangle>0\}$.
Suppose, for the contrary, that $\langle y-x,n\rangle\leq 0$. Parametrize the straight line segment $[x,y]$ linearly as $x(t),t\in[0,1]$, $x(0)=x$, $x(1)=y$. For each $t\in[0,1)$, let $u(t)=b^{\partial K}(x(t))$ be the only element $B^{\partial K}(x(t))$. Consider also $v(t)=c^{\partial K}(x(t))$, and notice that $v(1)=y$. The key idea is to look at the continuous family of straight line segments $[x(t),v(t)]$, and investigate how they intersect the open ball $O$ centered at $x$ of radius $r^{\partial K}(x)$. We make the following observations.
  • The open ball $O$ does not intersect $\mathcal{T}$, while $v(t)\in\mathcal{T}$. So for each $t\in[0,1)$, the segment $[x(t),v(t)]$ starts inside $O$ and ends outside of $O$.

### v2

@@ 6 Final Remarks
Although our results providing bounds on lifespans via widths are primarily theoretical, they offer practical value both in terms of improving the interpretability of PH features and in applications. Computing the exact Alexandrov, Urysohn, or Kolmogorov $k$-width for a given set $\mathcal{X}$ is often computationally challenging. However, by definition, these widths arise as infima of certain measurements over $k$-dimensional spaces (the “cores”). While identifying the optimal $k$-dimensional space could be highly complex, any meaningful and well-chosen $k$-dimensional space can yield relevant measurements that serve as upper bounds for these widths and, consequently, for the lifespans. Thus, even if calculating the optimal bound is infeasible, our results can be effectively leveraged to provide rough yet meaningful upper bounds for the lifespans of significant topological features. See 5.2 for other considerations related to potential uses of our bounds.
We highlight that our framework enables a bidirectional exchange of concepts, integrating ideas from metric geometry into applied algebraic topology and supporting their application in the reverse direction. Specifically, since lifespans are bounded above by various notions of width introduced in Section 3, they consequently provide lower bounds for these quantities.
In other words, for a given metric space $\mathcal{X}$, the maximum lifespan over classes in $\mathrm{Spec}_{k}(\mathcal{V}_{\bullet}(\mathcal{X}))$ (with zero birth time) serves as a lower bound for the corresponding width (e.g., $\operatorname{UW}_{k-1}(\mathcal{X})$).
This relationship offers a practical approach for estimating widths by leveraging topological persistence. We exemplify this through the following example.
Example 6.1 (The Urysohn Width of the $n$-Torus).
Let $n\geq 1$ be any integer and $a_{1}\geq a_{2}\geq\cdots\geq a_{n}>0$. Consider the $n$-torus $T^{n}:=a_{1}S^{1}\times a_{2}S^{1}\times\cdots\times a_{n}S^{1}$ endowed with the $\ell^{\infty}$ product metric, and where each $a_{i}S^{1}$ factor has the geodesic metric (with diameter $\pi a_{i}$). Then, we claim that for any $k\in\{1,\ldots,n\}$ we have
$\frac{2\pi}{3}a_{k}\leq\operatorname{UW}_{k-1}(T^{n})\leq\pi a_{k}.$
The upper bound can be trivially obtained by considering the projection onto the first $k-1$ factors of $T^{n}$. The lower bound can be obtained through an argument via PH as follows:
  1. (1)
    As proved by Adamaszek and Adams in [AA17], $I:=(0,\tfrac{2\pi}{3}]$ is the only bar in the VR-barcode of $S^{1}$ whose left endpoint is zero. Furthermore, this interval appears in degree-1.
  2. (2)
    By the Künneth formula for the VR-barcodes of $\ell^{\infty}$-products of compact metric spaces (see e.g. [LMO24, Theorem 6.1 and Example 6.4]), the only intervals in the degree-$k$ VR-barcode of $T^{n}$ with zero left endpoint must arise from intersecting exactly $k$ of the intervals $a_{1}I,a_{2}I,\cdots,a_{n}I$ each corresponding to the degree-1 VR-barcode of one of the $n$ $S^{1}$-factors of $T^{n}$. These intersections are precisely of the form $(0,d(L)]$ where $L\subset\{1,\ldots,n\}$ s.t. $|L|=k$ and
    $d(L):=\tfrac{2\pi}{3}\min_{\ell\in L}a_{\ell}.$
    Then, the maximum of $d(L)$ over all such subsets $L$ equals $\tfrac{2\pi}{3}a_{k}$.
  3. (3)
    By 3.3 and 2.34 we now conclude that $\tfrac{2\pi}{3}a_{k}\leq\operatorname{UW}_{k-1}(T^{n}).$
Compare with [Gro88, (E1)] and see also [Gro83, page 8].
Standard PCA vs. Lifespans.
@@
In Section 3.5, we give bounds for Čech lifespans via the variances $\{\nu_{k}\}$ induced by PCA (Section 1.2) As one can easily notice, in order to have such a rigorous bound, we modified the usual PCA definition and considered an $\ell^{\infty}$-variant. However, this makes the PCA structure highly sensitive to outliers in comparison with the original PCA. On the other hand, standard PCA is a mainstream, highly effective dimension reduction tool for real-life applications with several very efficient computational techniques available. While our results do not say anything about the relation between the original PCA and PH, an experimental result relating the lifespans of bars in the persistence diagram of the Čech filtration $\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbb{R}^{N})$ with variances $\widetilde{\nu}_{k+1}(\mathcal{X})$ of original PCA would be very interesting, for a given finite set of points $\mathcal{X}\subset\mathbb{R}^{N}$. It would be particularly interesting and useful for real-life applications to carry out a statistical comparative analysis for random finite subsets $\mathcal{X}$ in $\mathbb{R}^{N}$ (and for $k\geq 1$).
@@
One can notice that when describing the width-based arguments in Section 3, we first introduce a $k$-dimensional optimal core $\Lambda_{k}$ for a given set $\mathcal{X}$, then the $k$-width $\mathrm{W}_{k}(\mathcal{X})$ is defined as some kind of “distance” from $\mathcal{X}$ to $k$-core $\Lambda_{k}$. Hence, when $\mathrm{W}_{k}(\mathcal{X})$ is small, in metric geometry, $\mathcal{X}$ is regarded as “essentially $k$-dimensional”. There is a similar notion in statistics called Principal Curves and Surfaces; see Hastie and Stuetzle [HS89], Delicado [Del01], and Ozertem and Erdogmus [OE11]. While principal curves and surfaces are defined as $1$- and $2$-dimensional objects, one can easily generalize the idea to any dimension $k$, e.g. principal $k$-manifolds. In our setting, for a compact subset $\mathcal{X}$ in a Banach space $\mathbf{Y}$, principal curves and surfaces can be considered as $k$-dimensional objects $\Sigma_{k}$ which minimize the $\ell^{2}$-distance from $\mathcal{X}$ to $\Sigma_{k}$ for $k=1,2$ with some normalization condition on $\Sigma_{k}$. In this sense, principal curves and surfaces can be viewed as nonlinear generalizations of principal component analysis (PCA).
Similarly, for a given set $\mathcal{X}$, our $k$-cores and principal $k$-manifolds can be regarded as analogous constructs which extend the underlying idea to a broader geometric framework. In Section 3.5, we defined PCA as $\ell^{\infty}$-version of the original $\ell^{2}$ PCA . Similarly, our $k$-cores $\Lambda_{k}$ minimize the $\ell^{\infty}$-distance between $\mathcal{X}$ and $\Lambda_{k}$ while principal $k$-manifolds $\Sigma_{k}$ minimize the $\ell^{2}$-distance between $\mathcal{X}$ and $\Sigma_{k}$. By using this analogy, as principal curves and surfaces are suggested as dimension reduction method, one can consider our $k$-cores represent the essential structure of $\mathcal{X}$ when $\mathrm{W}_{k}(\mathcal{X})$ is small. Furthermore, just like the discussion in the preceding paragraph (PCA vs PCA), it would be interesting to carry out a statistical study of the relationship between the Čech lifespans and $\ell^{2}$-distance to principal $k$-manifolds.
@@ Appendix A Auxiliary Properties of Widths
Suppose $\operatorname{UW}_{k}(\mathcal{X})<\delta$, and let us show that $\operatorname{AW}_{k}(\mathcal{X}\subset\mathbf{Y})<\delta$. There is a continuous map $f:\mathcal{X}\to\Delta^{k}$ to a finite $k$-dimensional complex with fibers of diameter $<\delta$. Subdivide $\Delta^{k}$ very finely, so that the preimage of any open star of $\Delta^{k}$ under the map $f$ has diameter $<\delta$ [^16] Recall that the open star $S_{v}$ of a vertex $v\in\Delta^{k}$ is the union of the relative interiors of all simplices of $\Delta^{k}$ that contain $v$. For each vertex $v\in\Delta^{k}$, pick a point $c_{v}$ in the preimage of the open star $S_{v}$ (unless this preimage is empty; in this case, we can safely remove $S_{v}$ from $\Delta^{k}$). The ball $B_{\delta}(c_{v})\subset\mathbf{Y}$ of radius $\delta$ centered at $c_{v}$ covers $f^{-1}(S_{v})$. Consider an auxiliary map $\gamma:\Delta^{k}\to\mathbf{Y}$ defined by sending any vertex $v\in\Delta^{k}$ to $c_{v}\in\mathbf{Y}$, and then extending linearly on $\Delta^{k}$. Compose this map with $f$, and consider $\gamma\circ f:\mathcal{X}\to\mathbf{Y}$. Its image is a simplicial complex of dimension at most $k$, and to complete the proof it suffices to show that every point $x\in\mathcal{X}$ is moved by distance $\|x-\gamma(f(x))\|<\delta$. Let $f(x)$ lie in the relative interior of a simplex of $\Delta^{k}$ with the vertices $v_{0},\ldots,v_{m}$. Then $x\in f^{-1}(S_{v_{i}})\subset B_{\delta}(c_{v_{i}})$, and $\|x-c_{v_{i}}\|<\delta$, where $0\leq i\leq m$. By construction of $\gamma$, the point $\phi(f(x))$ lies in the convex hull of the points $c_{v_{i}}$, $0\leq i\leq m$. Therefore, $\|x-\gamma(f(x))\|$ does not exceed the maximum of $\|x-c_{v_{i}}\|$ over $0\leq i\leq m$, and this maximum is less than $\delta$. ∎
@@
Denote $\mathcal{X}=\bigcap\mathcal{X}_{i}$, $w=\operatorname{TW}_{k}(\mathcal{X}\subset\mathbf{Y})$, and let $f:\mathcal{X}\to\mathbf{Y}$ be a witness map: a continuous map whose image lies in a finite contractible $k$-dimensional simplicial complex $\Delta\subset Y$, and such that $\|x-f(x)\|<w+\epsilon/3$ for all $x\in\mathcal{X}$ and some arbitrarily chosen $\epsilon>0$. For each point $y\in\Delta$, the fiber $f^{-1}(y)$ lies in the open ball $B_{w+\epsilon/3}(y)$ (note that $f^{-1}(y)$ can be empty). Since $f$ is continuous on the compact set $\mathcal{X}$, it is uniformly continuous, hence there is a tiny radius $\rho>0$ such that such that the “thickened fiber” $f^{-1}(B_{\rho}(y))$ lies in the open ball $B_{w+2\epsilon/3}(y)$, for each $y\in\Delta$. Since $\Delta$ is a finite complex of dimension $k$, there is a finite closed cover $\bigcup D_{j}=\Delta$ such that its multiplicity is $k+1$, its nerve is homotopy equivalent to $\Delta$, each $D_{j}$ fits in a ball $B_{\rho}(y_{j})$, and $C_{j}:=f^{-1}(D_{j})$ fits in the ball $B_{w+2\epsilon/3}(y_{j})$.
@@
is attained and positive. Take a positive $\delta$ smaller than $\epsilon/3$ and also smaller than each $\delta_{J}$ over all collections $J$ such that $\bigcap\limits_{j\in J}C_{j}=\varnothing$. Consider the open cover $\{U_{j}\}$ of $\mathcal{X}$, where $U_{j}=\mathcal{N}_{\delta}(C_{j}\subset\mathbf{Y})$. It has the same nerve as $\{C_{i}\}$, and each $U_{j}$ is contained in $B_{w+\epsilon}(y_{j})$. We have
$\mathcal{X}\subset\bigcup_{j}U_{j}=\mathcal{N}_{\delta}(\mathcal{X}\subset\mathbf{Y}).$
Now pick a partition of unity $\{\psi_{j}\}$ subordinate to $\{U_{j}\}$. Use it to map $\bigcup_{j}U_{j}$ to its nerve. Namely, a point $x\in\bigcup_{j}U_{j}$ is mapped to $\sum\limits_{j}\psi_{j}(x)y_{j}$. This gives a (possibly non-surjective) map $\bigcup_{j}U_{j}\to\Delta$. Since $U_{j}\subset B_{w+\epsilon}(y_{j})$, every point under this map is shifted by distance less than $w+\epsilon$.
@@
Here we only work in $\mathbf{Y}=\mathbb{R}^{N}$ with Euclidean metric, and we are interested in 00-übercontractible sets, which we call übercontractible. A rich source of übercontractible sets comes from the cut-locus construction, usually considered in a more general Riemannian setting; see, e.g., Wolter [Wol85]. A very close concept is that of skeleta (see, e.g., Fremlin [Fre97]), different from our definition only in that we take the closure. In computational geometry, low-dimensional skeleton construction are also called medial axes; see, e.g., the survey by Saha, Borgefors and Sanniti di Baja [SBdB16]. We only formulate the definition of the cut-locus for convex hypersurfaces in $\mathbf{Y}=\mathbb{R}^{N}$.
@@
The cut-locus $C^{\partial K}$ of the boundary of a convex polytope $K$ is übercontractible.
The key lemma in the proof comes from the Morse theory for the distance functional. It was applied by Bárány, Holmsen and Karasev in [BHK15, Section 3] to give a sufficient condition for a set to be contractible, but the proof there, in fact, guarantees that the set is übercontractible.
Lemma B.3 (cf. [BHK15, Theorem 2]).
Let $\mathcal{T}\subset\mathbf{Y}$ be a union of finitely many compact convex sets. For each $x\notin\mathcal{T}$ consider the largest ball centered at $x$ whose interior does not meet $\mathcal{T}$. Let $B^{\mathcal{T}}(x)$ be the set of the points where this ball touches $\mathcal{T}$. Suppose that for each $x\notin\mathcal{T}$ we have $x\notin\mathrm{conv}(B^{\mathcal{T}}(x))$. Then $\mathcal{T}$ is übercontractible.
Proof of B.2.
@@
We conclude this section by speculating how B.2 can be proven for convex bodies other than polytopes. Unfortunately, the case of a general convex body $K$ cannot be proven by approximating $K$ with polytopes $K_{i}$, because in order to guarantee $C^{0}$-convergence $C^{\partial K_{i}}\to C^{\partial K}$, we cannot get away with just polytopes. We sketch an argument that works for fairly general convex bodies, modulo some technicalities. We take a different approach and assume the following “tameness” assumption: $C^{\partial K}$ consists of finitely many compact convex sets. It allows us to apply B.3, but it is highly likely that the lemma can also be stated and proven in greater generality. Not only the class of tame convex bodies includes polytopes, but it also seems to be $C^{2}$-dense among all $C^{2}$-smooth convex bodies (we do not discuss this in detail since this digresses too far from the main topic of the paper).
Sketch of the proof of B.2 for tame convex bodies.
The following two properties will be shown to imply the übercontractibility of $C^{\partial K}$.
@@
    The cut-locus of $\partial K$ consists of finitely many compact convex sets (so we can use B.3).
@@
    The boundary of $K$ is $C^{2}$-smooth and strongly convex in the sense that its second fundamental form is positive definite.
The tameness assumption guarantees the first property. We argue that we can inflate $K$ slightly to get a new convex body $K^{\prime}$ that satisfies the second property, while preserving cut-locus $C^{\partial K^{\prime}}=C^{\partial K}=\mathcal{T}$. This amounts to choosing $r^{\partial K^{\prime}}(x)$ for $x\in\mathcal{T}$ carefully, so that on every face of $\mathcal{T}$, $r^{\partial K^{\prime}}$ is $C^{1}$-close to $r^{\partial K}$, and $r^{\partial K^{\prime}}$ is strongly convex. The details are omitted.
In the rest of the proof we assume that $K$ satisfies the two properties above, and we will verify the assumption of B.3 for $\mathcal{T}=C^{\partial K}$ to show that it is übercontractible. Pick points $x\in K\setminus\mathcal{T}$, and $y\in B^{\mathcal{T}}(x)$. Our goal is to prove that $\langle y-x,n\rangle>0$, where $n$ is the inner normal to $\partial K$ at $b^{\partial K}(x)$. This will imply that $x\notin\mathrm{conv}(B^{\mathcal{T}}(x))\subset\{z\in\mathbf{Y}:\langle z-x,n\rangle>0\}$.
Suppose, for the contrary, that $\langle y-x,n\rangle\leq 0$. Parametrize the straight line segment $[x,y]$ linearly as $x(t),t\in[0,1]$, $x(0)=x$, $x(1)=y$. For each $t\in[0,1)$, let $u(t)=b^{\partial K}(x(t))$ be the only element $B^{\partial K}(x(t))$. Consider also $v(t)=c^{\partial K}(x(t))$, for $t\in[0,1]$, and notice that $v(1)=y$. The key idea is to look at the continuous family of straight line segments $[x(t),v(t)]$, and investigate how they intersect the open ball $O$ centered at $x$ of radius $r^{\partial K}(x)$. We make the following observations.
  • The open ball $O$ does not intersect $\mathcal{T}$, and $v(t)\in\mathcal{T}$. So for each $t\in[0,1)$, the segment $[x(t),v(t)]$ starts inside $O$ and ends outside of $O$.